OpenStudy (anonymous):

Which factor is canceled from the numerator and denominator to simplify x^2 - x - 12 / x^2 - 16?

OpenStudy (anonymous):

\[\frac{x^2-x-12}{x^2-16}=\frac{(x-4)(x+3)}{(x-4)(x+4)}\]